Consumer buys 10 units of Good A when the price of Good B is $5. When the price of Good B rises to $6 (the price of Good A remaining unchanged) the consumer buys 14 units of Good A.
Part A (6 MARKS)
Using an appropriate formula, calculate this Consumer’s cross Elasticity of demand for Good A. Show your working.
Part B (4 MARKS)
Is Good A, a substitute for, or a complement to, Good B? Explain your reasoning.
1
Expert's answer
2021-08-31T16:21:33-0400

cross elasticity of demand is given by

"(Change Qa \u00f7 change Pb) \u00f7 (Pb \u00f7Qa)"

change in quantity A = 14 - 10 = 4 units

change in price B = 6 -5 = 1

therefore "CES = 4 * 6 \u00f7 14"

=1.71


B. Good A is a substitute to B since it has a positive cross elasticity of demand